sheila practiced 1 1/2 hours monday, 3 3/4 hours tueday, and 2 1/4 hours wednesday. what is the toatl number of hours that she practiced?
Found 2 solutions by RAY100, nerdybill:
Answer by RAY100(1637)

M= 1 1/2 = 1 2/4
.
T=,,,,,,,,,3 3/4
.
W=,,,,,,,,,2 1/4
.
TOTAL = ,,6 + 6/4 = 6+1+2/4 =7 +2/4 = 7+1/2 = 7 1/2 hrs
.
check,,,1.5 + 3.75 + 2.25 = 7.5,,,,ok
.

Answer by nerdybill(5411)

sheila practiced 1 1/2 hours monday, 3 3/4 hours tueday, and 2 1/4 hours wednesday. what is the toatl number of hours that she practiced?
.
1 1/2 + 3 3/4 + 1 1/4
= (1 + 3 + 1) + (1/2 + 3/4 + 1/4)
= (5) + (2/4 + 3/4 + 1/4)
= 5 + (2+3+1)/4
= 5 + 6/4
= 5 + 3/2
= 5 + 1 1/2
= 6 1/2 hours